OpenStudy (anonymous):

Which Property of Equality would be best to use first to solve the equation -9n + 5 = -67 for n? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

idk what it's called but i do know how to solve for n. xP

OpenStudy (anonymous):

if i show you how, maybe you could look for the definition of each property. -9n + 5 = -6 1st.. transfer -9n to right side. 5 = -6 + 9n 2nd.. transfer -6 to left side. 5 + 6 = 9n 3rd.. divide both sides by 9 to eliminate 9 in 9n.. 11/9 = n or n = 11/9

OpenStudy (unklerhaukus):

hbaldon's 1st.. transfer, uses the fact that we can add something to both sides of the equation and this will have the same solution as the solution to the original hbaldron has added 9n to both sides